In a triangle, one of the angles is 100 degrees, the measure of the middle angle is 24 degrees more than the measure of the smallest angle. What are the measure of these two angles
Answer by colliefan(242)

The sum of all angles in a triangle is always 360.
One angle can be represented by x.
The other is then x+24.
If the other angle is 100, then these two must equal 260 when added together.
x + (x+24)= 260
x + x + 24= 260
2x + 24= 260
2x + 24 - 24= 260 - 24
2x + 0= 236
2x = 236
2x*1/2 = 236*1/2
x*2*1/2 = 236*1/2
x*1 = 118
x = 118
If x or 118 deg is one angle, the other is x+24 or 142 deg.






You could also start with
x + (x+24) + 100 = 360
and you would get the same result.